# Scene Coordinate Reconstruction: Posing of Image Collections via Incremental Learning of a Relocalizer

### v1

@@ Abstract
We address the task of estimating camera parameters from a set of images depicting a scene. Popular feature-based structure-from-motion (SfM) tools solve this task by incremental reconstruction: they repeat triangulation of sparse 3D points and registration of more camera views to the sparse point cloud. We re-interpret incremental structure-from-motion as an iterated application and refinement of a visual relocalizer, that is, of a method that registers new views to the current state of the reconstruction. This perspective allows us to investigate alternative visual relocalizers that are not rooted in local feature matching. We show that scene coordinate regression, a learning-based relocalization approach, allows us to build implicit, neural scene representations from unposed images. Different from other learning-based reconstruction methods, we do not require pose priors nor sequential inputs, and we optimize efficiently over thousands of images. Our method, ACE0, estimates camera poses to an accuracy comparable to feature-based SfM, as demonstrated by novel view synthesis.
@@ 1 Introduction
The success of Neural Radiance Fields (NeRF) [61] has renewed interest in the question of whether SfM can be solved differently, based on neural, implicit scene representations rather than 3D point clouds. There has been some progress in recent years but, thus far, learning-based approaches to camera pose recovery still have significant limitations. They either require coarse initial poses [43, 100, 53, 103], prior knowledge of the pose distribution [60] or roughly sequential inputs [10]. In terms of the number of images that learning-based approaches can handle, they are either explicitly targeted at few-frame problems [97, 101, 84, 52, 105, 98] or they are computationally so demanding that they can realistically only be applied to a few hundred images at most [100, 53, 10]. We show that none of these limitations are an inherent consequence of using learning-based scene representations.
@@
We show that “scene coordinate regression”, an approach to visual relocalization proposed a decade ago [83], has the desirable properties and can serve as the core of a new approach to learning-based SfM: Scene coordinate reconstruction. Rather than optimizing over image-to-image matches, like feature-based SfM, scene coordinate reconstruction regresses image-to-scene correspondences directly. Rather than representing the scene as a 3D point cloud with high dimensional descriptors, we encode the scene into a lightweight neural network. Our approach works on unsorted images without pose priors and efficiently optimises over thousands of images, see Fig. 1.
@@ 2 Related Work
SIFT [58] and other robust feature descriptors allow matching image features across wide baselines enabling systems to reconstruct large-scale scenes using Internet images [86, 87, 102, 82]. Feature tracks are built by matching the local features between two or more images. In our work we do not explicitly compute matches between images, which can be computationally expensive.
Feature tracks and estimated relative poses are used to solve for the 3D feature coordinates, camera poses and calibrations (intrinsic matrices). This geometric optimization problem is mainly solved using bundle adjustment which was explored in photogrammetry and geodesy [18, 49] and became standard in the computer vision community [94, 39]. Bundle adjustment relies on the initialization being close to the solution (i.e., camera poses and 3D points are already mostly accurate). There are two main approaches to this problem. Incremental SfM [6, 68] starts the reconstruction from very few images to create a high-quality seed reconstruction that progressively grows by registering more images and refining the reconstruction until convergence. Global SfM approaches solve for “global” poses of all images using estimates of relative poses, i.e., motion averaging [36, 37], rotation averaging [38, 59] and pose-graph optimization [20]. Our work is similar to Incremental SfM, as we also progressively register images to the reconstructed scene starting from a seed reconstruction.
Various techniques were proposed to improve SfM runtime for very large sets of images [27, 2, 1, 40, 8, 9, 88, 35, 93].
@@
Some learning-based approaches encode the map of a scene in the weights of a neural network. PoseNet [46] directly regresses the absolute camera pose given an input image using a CNN that was trained on image-pose pairs of the reconstructed scene. Sattler et al. show that extrapolating outside the mapped poses can be challenging for absolute pose regression approaches [79]. Relative pose regression networks [4, 31, 50, 3, 95] estimate the relative pose for a pair of images, allowing triangulation of the query image from multiple map images, or even a single-image. Scene coordinate regression [83, 22, 13, 14, 21, 16, 11, 51] directly predicts 3D coordinates of a point given a patch. This approach generalizes well as the camera pose is solved using PnP [34] within a RANSAC loop [33]. ACE [11] shows that training of scene coordinate regression can be greatly accelerated. In our work we train the ACE localizer, but we start from images without poses.
@@ 3 Method
We also want to recover the 3D structure of the scene: Each pixel $j$ in image $i$ with 2D pixel position $\mathbf{p}_{ij}$ has a corresponding coordinate in 3D, denoted as $\mathbf{y}_{ij}$. 2D pixel positions and 3D scene coordinates are related by the camera pose and the projection function $\boldsymbol{\pi}$:
@@ 3.1 Neural Mapping
We train the scene coordinate regression model iteratively where we denote the current time step as $t$ and the corresponding scene model as $f^{t}_{\text{SCR}}$. We iterate between training the scene model, and registering new views, see Fig. 2.
@@
Mapping Framework. We optimize Eq. 3 using stochastic gradient descent, using the fast-learning scene coordinate regressor ACE [11] (Accelerated Coordinate Encoding). ACE is trained in minutes, even for thousands of views. Since we have to train the model in multiple iterations, each individual iteration must be done quickly. ACE approximates Eq. 3 by sampling random patches from the training set, $\mathcal{I}^{t}_{\text{Reg}}$. To do that efficiently, ACE employs a pre-trained encoder to pre-compute high dimensional features for a large number of training patches. The encoder stays frozen during mapping. The actual scene model, which is trained, is a multi-layer perceptron (MLP) that maps encoder features to scene coordinates; see Fig. 2 (top right) for a visual representation.
@@
Pose Refinement. Differently from ACE [11] protocol, the ground truth poses $T^{GT}_{j}$ are unknown during training. Instead, we have $T^{t}_{j}$, estimates based on earlier iterations of the reconstruction. Since these estimates can be inaccurate, we add the ability to refine poses during mapping. We implement refinement using an MLP:
$T_{j}^{t}=f^{t}_{\text{Pose}}(\tilde{T}^{t}_{j}),$ (4)
where $\tilde{T}^{t}_{j}$ denotes the initial pose estimate at the start of a mapping iteration. The refinement MLP ingests $\tilde{T}^{t}_{j}$ as $3\times 4=12$ values and predicts 12 additive offsets. We jointly optimize $f^{t}_{\text{Pose}}$ and $f^{t}_{\text{SCR}}$ to minimize reprojection error 3. We orthonormalize the rotation using Procrustes [17] to make the pose a valid rigid body transformation. We show the impact of pose refinement for one iteration in Fig. 3 (left).
We discard the refinement MLP after each mapping iteration. Its purpose is to enable the scene model, $f^{t}_{\text{SCR}}$, to converge to a consistent scene representation. With neither poses nor 2D-3D correspondences fixed in Eq. 3, the scene coordinate regressor could drift or degenerate. As regularization, we optimize the pose refiner $f^{t}_{\text{Pose}}$ using AdamW [57] with weight decay, biasing the MLP to predict small updates. We found no further regularization necessary. The smoothness prior of the networks [96] encourages a multi-view consistent solution.
As an alternative, we could back-propagate directly to the input poses [104, 100] to update them. However, when optimizing over thousands of views, the signal for each single pose becomes extremely sparse. Cameras are also correlated via the scene representation. If the optimization removes drift in the scene representation, multiple cameras have to be moved. The refinement MLP is an effective means to model the correlation of cameras.
Calibration Refinement. We do not assume information about precise calibration parameters, although that information is often reported by devices and, e.g. stored in the EXIF data. We do rely on a sensible initialization, assuming that the principal point is in the center, that pixels are unskewed and square. We do not model image distortion. While these are reasonable assumptions for many data regimes, we cannot rely on the focal length to be given. Thus, we refine the focal length via a utility function starting from a heuristic: $K_{i}^{t}=f^{t}_{\text{Calib}}(K^{\text{init}}_{i})$. The refinement function entails a single learnable parameter $\alpha^{t}$ such that the focal length $f^{t}_{i}=f^{\text{init}}\cdot(1+\alpha^{t})$. We optimize $\alpha^{t}$ using AdamW [57] with weight decay, biasing it towards a small relative scale factor. We set $f^{\text{init}}$ to 70% of the image diagonal and it is typically a parameter shared by all cameras.
@@ 3.2 Relocalization
$\tilde{T}^{t+1}_{i},s^{t+1}_{i}=g(K^{t}_{i},\{(\mathbf{p}_{ij},\mathbf{y}^{t}_{ij})\}).$ (5)
Here, we assume that the pose solver returns a confidence score $s^{t+1}_{i}$ alongside the pose itself that lets us decide whether the pose of the image has been estimated successfully. We simply utilize the inlier count as score $s^{t+1}_{i}$ and apply a threshold to form the training set of the next mapping iteration: $\mathcal{I}^{t+1}_{\text{Reg}}=\{I_{i}|s^{t+1}_{i}>\tau_{s}\}$. The relocalization process is depicted in Fig. 2, bottom.
@@ 3.3 Initialization
We start the reconstruction with one image: $\mathcal{I}^{0}_{\text{Reg}}=\{I_{\text{seed}}\}$. We set the seed pose $T^{0}_{\text{seed}}$ to identity, and we initialize the calibration $K^{\text{init}}_{\text{seed}}$ as explained above.
We cannot train a scene coordinate regression network using Eq. 3 with a single image. The reprojection error is ill-conditioned without multiple views constraining the depth. Therefore, we optimize a different objective in the seed iteration, inspired by Map-free Relocalization [3]. Arnold et al. argue that a single image and a depth estimate are sufficient to relocalize query images. Our experiments show that a few coarse estimates are enough to serve as initialization for optimizing the reprojection error according to Eq. 3.
@@ 3.4 Implementation
We report more implementation details and hyper-parameters in the supplement. We will also make our code publicly available to ensure reproducibility.
@@ 4 Experiments
Baselines. We consider the pose estimates of COLMAP with default parameters as our pseudo ground truth, obtained by extensive processing. We also run COLMAP with parameter settings recommended for large image collections of 1k images and more [81]. This variation achieves much faster processing for a slight drop in quality (denoted COLMAP fast). We use a V100 GPU for COLMAP feature extraction and matching, and we specify the exact parameters of COLMAP in the supplement. Furthermore, we show some results of RealityCapture [70], an efficient commercial feature-based SfM pipeline.
@@
Benchmark. We show results on 7-Scenes [83], a relocalization dataset, on Mip-NeRF 360 [5], a view synthesis dataset and on Tanks and Temples [48], a reconstruction dataset. Comparing poses on these datasets is problematic, as our pseudo ground truth is estimated rather than measured. For example, an approach might be more accurate than COLMAP on individual scenes. Computing pose errors w.r.t. COLMAP would result in incorrect conclusions. Therefore, we gauge the pose quality in a self-supervised way, using novel view synthesis.
We let each method estimate the poses of all images of a scene. For evaluation, we split the images into training and test sets. We train a Nerfacto [91] model on the training set, and synthesize views for the test poses. We compare the synthesized images to the test images and report the difference as peak signal-to-noise ratio (PSNR). To ensure a fair comparison to NeRF-based competitors, we use these methods in the same way as the other SfM methods: we run them to pose all images, and train a Nerfacto model on top of their poses. This is to ensure that we solely compare the pose quality across methods, and not different capabilities in view synthesis. The quality of poses affects the PSNR numbers in two ways: Good training poses let the NeRF model fit a consistent scene representation. Good testing poses make sure that the synthesized images are aligned with the original image. We explain further details including perceptual metrics in the supplement. In spirit, our evaluation is similar to the Tanks and Temples benchmark which evaluates a derived scene mesh rather than camera poses. However, our evaluation can be applied to arbitrary datasets as it does not need ground truth geometry.
@@
 Pseudo Ground Truth  All Frames  200 Frames  50 Frames Frames  Kinect  COLMAP  COLMAP  DROID-SLAM  ACE0  KF+ACE0  BARF  NoPE-NeRF  ACE0  DUSt3R  ACE0 Fusion (default) (fast) [92] (ours) (ours) [53] [10] (ours) [98] (Ours) Chess 6k 19.6 23.8  23.5  19.3  23.3  23.0 12.8 12.6 20.9 18.9 19.2 Fire 4k 19.2 22.3  22.4  13.0  22.3  22.3 12.7 11.8 21.3 18.8 19.5 Heads 2k 17.0 19.1  18.9  17.6  18.8  19.1 10.7 11.8 19.6 18.4 21.3 Office 10k 18.9 21.5  21.3 failed  21.1  21.5 11.9 10.9 19.7 12.5 13.7 Pumpkin 6k 19.9 24.0  23.8  18.3  24.0  23.8 19.6 14.2 24.2 21.7 22.3 RedKitchen 12k 17.6 21.4  21.4  10.9  20.8  20.9 11.6 11.2 18.5 13.8 13.7 Stairs 3k 19.0 21.0  15.8  13.0  17.7  19.9 15.8 15.9 18.0 15.3 15.4 Average 18.7 21.9  21.0 N/A  21.2  21.5 13.6 12.6 20.3 17.1 17.9 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
Discussion. We show results in Table 1. Of both pseudo ground truth versions, KinectFusion achieves lower PSNR numbers than COLMAP, presumably due to the lack of global optimization. COLMAP with fast parameters shows only slightly lower PSNR numbers than COLMAP with default parameters. Both versions of running COLMAP take considerable time to reconstruct each scene. We note that COLMAP has been optimised for quality, rather than speed. Not all acceleration strategies from the feature-based SfM literature have been implemented in COLMAP, so presumably comparable quality can be obtained faster. DROID-SLAM [92] largely fails on 7-Scenes, presumably due to the jumps between individual scans of each scene.
Our approach, ACE0, achieves a pose quality comparable to the COLMAP pseudo ground truth while reconstructing each scene in $\sim$1 hour despite the large number of images. We show qualitative examples in Figure 4. We also demonstrate that ACE0 can swiftly optimize an initial set of approximate poses. When starting from KinectFusion poses, ACE0 increases PSNR significantly in less than 10 minutes per scene. In the supplement, we include a parameter study on 7-Scenes to show that ACE0 is robust to the choice of depth estimator. We also show the positive impact of pose refinement on the reconstruction quality as well as the reconstruction speedup due to our early stopping schedule.
For our learning-based competitors, we sub-sampled images due to their computational constrains, see right side of Table 1. Even using only 200 images, NoPe-NeRF [10] takes 2 days to fit a model and estimate poses. Despite these long processing times, we observe poor pose quality of BARF and NoPe-NeRF in terms of PSNR. BARF [53] requires pose initialisation. We provide identity poses since the scenes of 7Scenes are roughly forward-facing. Still, the camera motions are too complex for BARF to handle. NoPe-NeRF does not require pose initialisation but needs roughly sequential images, which we did provide. NoPe-NeRF relies upon successive images having similar poses to perform well, and struggles with the large jumps between the subsampled images.

### v2

@@ Abstract
We address the task of estimating camera parameters from a set of images depicting a scene. Popular feature-based structure-from-motion (SfM) tools solve this task by incremental reconstruction: they repeat triangulation of sparse 3D points and registration of more camera views to the sparse point cloud. We re-interpret incremental structure-from-motion as an iterated application and refinement of a visual relocalizer, that is, of a method that registers new views to the current state of the reconstruction. This perspective allows us to investigate alternative visual relocalizers that are not rooted in local feature matching. We show that scene coordinate regression, a learning-based relocalization approach, allows us to build implicit, neural scene representations from unposed images. Different from other learning-based reconstruction methods, we do not require pose priors nor sequential inputs, and we optimize efficiently over thousands of images. In many cases, our method, ACE0, estimates camera poses with an accuracy close to feature-based SfM, as demonstrated by novel view synthesis.
@@ 1 Introduction
The success of Neural Radiance Fields (NeRF) [63] has renewed interest in the question of whether SfM can be solved differently, based on neural, implicit scene representations rather than 3D point clouds. There has been some progress in recent years but, thus far, learning-based approaches to camera pose recovery still have significant limitations. They either require coarse initial poses [45, 103, 55, 106], prior knowledge of the pose distribution [62] or sequential inputs [11, 94, 10]. In terms of the number of images that learning-based approaches can handle, they are either explicitly targeted at few-frame problems [99, 104, 86, 54, 108, 101] or they are computationally so demanding that they can realistically only be applied to a few hundred images at most [103, 55, 10]. We show that none of these limitations are an inherent consequence of using learning-based scene representations.
@@
We show that “scene coordinate regression”, an approach to visual relocalization proposed a decade ago [85], has the desirable properties and can serve as the core of a new approach to learning-based SfM: Scene coordinate reconstruction. Rather than optimizing over image-to-image matches, like feature-based SfM, scene coordinate reconstruction regresses image-to-scene correspondences directly. Rather than representing the scene as a 3D point cloud with high dimensional descriptors, we encode the scene into a lightweight neural network. Our approach works on unsorted images without pose priors and efficiently optimises over thousands of images, see Figure 1.
@@ 2 Related Work
SIFT [60] and other robust descriptors allow matching image features across wide baselines enabling systems to reconstruct large-scale scenes using Internet images [88, 89, 105, 84]. Image-to-image matches can also be regressed directly in a detector-free setup [91, 41]. Feature tracks across multiple images are built from image-to-image matches. Feature tracks and estimated relative poses are used to solve for the 3D feature coordinates, camera poses and calibrations (intrinsic matrices). This geometric optimization problem is mainly solved using bundle adjustment which was explored in photogrammetry and geodesy [19, 51] and became standard in the computer vision community [96, 40]. Bundle adjustment relies on the initialization being close to the solution (i.e., camera poses and 3D points are already mostly accurate).
There are two main approaches to this problem. Incremental SfM [6, 70] starts the reconstruction from very few images to create a high-quality seed reconstruction that progressively grows by registering more images and refining the reconstruction until convergence. Global SfM approaches solve for “global” poses of all images using estimates of relative poses, i.e., motion averaging [37, 38], rotation averaging [39, 61] and pose-graph optimization [21]. Various techniques were proposed to improve SfM runtime for very large sets of images [28, 2, 1, 42, 8, 9, 90, 36, 95]. Our work is similar to Incremental SfM, as we also progressively register images to the reconstructed scene starting from a seed reconstruction. However, we do not explicitly compute image matches, nor feature tracks across images, which can be computationally expensive.
@@
Some learning-based approaches encode the map of a scene in the weights of a neural network. PoseNet [48] directly regresses the absolute camera pose given an input image using a CNN that was trained on image-pose pairs of the reconstructed scene. Sattler et al. show that extrapolating outside the mapped poses can be challenging for absolute pose regression approaches [81]. Relative pose regression networks [4, 32, 52, 97] estimate the relative pose for a pair of images, allowing triangulation of the query image from multiple map images, or estimate a scale-metric relative pose w.r.t. a single map image [3]. Scene coordinate regression [85, 23, 14, 15, 22, 17, 12, 53] directly predicts 3D coordinates of a point given a patch. This approach generalizes well as the camera pose is solved using PnP [35] within a RANSAC loop [34]. ACE [12] shows that training of scene coordinate regression can be greatly accelerated. In our work we train the ACE localizer, but we start from images without poses.
@@ 3 Method
We also want to recover the 3D structure of the scene: Each pixel $j$ in image $i$ with 2D pixel position $\mathbf{p}_{ij}$ has a corresponding coordinate in 3D, denoted as $\mathbf{y}_{ij}$. The 2D pixel positions and 3D scene coordinates are related by the camera pose and the projection function $\boldsymbol{\pi}$:
@@ 3.1 Neural Mapping
We train the scene coordinate regression model iteratively where we denote the current time step as $t$ and the corresponding scene model as $f^{t}_{\text{SCR}}$. We iterate between training the scene model, and registering new views, see Figure 2.
@@
Mapping Framework. We optimize Eq. 3 using stochastic gradient descent, using the fast-learning scene coordinate regressor ACE [12] (Accelerated Coordinate Encoding). ACE is trained in minutes, even for thousands of views. The training speed is important since we have to train the model in multiple iterations. ACE approximates Eq. 3 by sampling random patches from the training set, $\mathcal{I}^{t}_{\text{Reg}}$. To do that efficiently, ACE employs a pre-trained encoder to pre-compute high dimensional features for a large number of training patches. The encoder stays frozen during mapping. The actual scene model, which is trained, is a multi-layer perceptron (MLP) that maps encoder features to scene coordinates; see Fig. 2 (top right) for a visual representation.
@@
Pose Refinement. Differently from the ACE [12] protocol, the ground truth poses $T^{\text{GT}}_{i}$ are unknown during training. Instead, we have $T^{t}_{i}$, estimates based on earlier iterations of the reconstruction. Since these estimates can be inaccurate, we add the ability to refine poses during mapping. We implement refinement using an MLP:
$T_{i}^{t}=f^{t}_{\text{Pose}}(\tilde{T}^{t}_{i}),$ (4)
where $\tilde{T}^{t}_{i}$ denotes the initial pose estimate at the start of a mapping iteration. Inspired by [112], the refinement MLP ingests $\tilde{T}^{t}_{i}$ as $3\times 4=12$ values and predicts 12 additive offsets. We orthonormalize rotations using Gram-Schmidt [18]. We jointly optimize $f^{t}_{\text{Pose}}$ and $f^{t}_{\text{SCR}}$ to minimize the reprojection error of Eq. 3. We show the impact of pose refinement for one iteration in Fig. 3 (left). We discard the refinement MLP after each mapping iteration. Its purpose is to enable the scene model, $f^{t}_{\text{SCR}}$, to converge to a consistent scene representation.
With neither poses nor 2D-3D correspondences fixed in Eq. 3, the scene coordinate regressor could drift or degenerate. As regularization, we optimize the pose refiner $f^{t}_{\text{Pose}}$ using AdamW [59] with weight decay, biasing the MLP to predict small updates relative to the initial estimates $\tilde{T}^{t}_{i}$. This relies on the assumption that mapping images in $\mathcal{I}^{t}_{\text{Reg}}$ have been registered close to their true position. If that assumption holds, the smoothness prior of the networks [98] encourages a multi-view consistent solution as shown in previous work [15, 17, 12].
As an alternative to an MLP refiner, we could back-propagate directly to the input poses [107, 103]. However, when optimizing over thousands of views, the signal for each single pose becomes sparse. Cameras are correlated via the scene representation. If the optimization removes drift in the scene, multiple cameras need to move. The MLP refiner models the correlation of cameras.
Calibration Refinement. We do not assume information about precise calibration parameters, although often reported by devices. We do assume that the principal point is in the center, that pixels are unskewed and square. We do not model image distortion. While these are reasonable assumptions for many data regimes, we cannot rely on the focal length to be given. Thus, we refine the focal length starting from a heuristic: $K_{i}^{t}=f^{t}_{\text{Calib}}(K^{\text{init}}_{i})$. The refinement function $f^{t}_{\text{Calib}}$ entails a single learnable parameter $\alpha^{t}$ such that the focal length $f^{t}_{i}=f^{\text{init}}\cdot(1+\alpha^{t})$. As before, superscript $t$ denotes the time step. We optimize $\alpha^{t}$ using AdamW [59] with weight decay, biasing it towards a small relative scale factor. Estimates of $\alpha$ are carried over across iterations. We set $f^{\text{init}}$ to 70% of the image diagonal and it is shared by all cameras in our experiments.
@@ 3.2 Relocalization
$\tilde{T}^{t+1}_{i},s^{t+1}_{i}=g\left(K^{t}_{i},\left\{\left(\mathbf{p}_{ij},\mathbf{y}^{t}_{ij}\right)\right\}\right).$ (5)
Here, we assume that the pose solver returns a confidence score $s^{t+1}_{i}$ alongside the pose itself that lets us decide whether the pose of the image has been estimated successfully. We simply utilize the inlier count as score $s^{t+1}_{i}$ and apply a threshold to form the training set of the next mapping iteration: $\mathcal{I}^{t+1}_{\text{Reg}}=\left\{I_{i}|s^{t+1}_{i}>\tau_{s}\right\}$. The relocalization process is depicted in Figure 2, bottom.
@@ 3.3 Initialization
We start the reconstruction with one image: $\mathcal{I}^{0}_{\text{Reg}}=\left\{I_{\text{seed}}\right\}$. We set the seed pose $T^{0}_{\text{seed}}$ to identity, and we initialize the calibration $K^{\text{init}}_{\text{seed}}$ as explained above.
We cannot train a scene coordinate regression network using Eq. 3 with a single image. The reprojection error is ill-conditioned without multiple views constraining the depth. Therefore, we optimize a different objective in the seed iteration, inspired by Map-free Relocalization [3]. Arnold et al. argue that a single image and a depth estimate allow to relocalize query images, albeit with limited accuracy. Our experiments show that such coarse relocalizations of a few images suffice as initialization for optimizing the reprojection error of Eq. 3.
@@ 3.4 Implementation
We report more implementation details and hyper-parameters in the supplement. Our code is also publicly available to ensure reproducibility.
@@ 4 Experiments
Baselines. We consider the pose estimates of COLMAP with default parameters as our pseudo ground truth, obtained by extensive processing. We also run COLMAP with parameter settings recommended for large image collections of 1k images and more [83]. This variation achieves much faster processing (denoted COLMAP fast). We use a V100 GPU for COLMAP feature extraction and matching, and we specify the exact parameters of COLMAP in the supplement. Furthermore, we show some results of RealityCapture [72], an efficient commercial feature-based SfM pipeline.
@@
Benchmark. We show results on 7-Scenes [85], a relocalization dataset, on Mip-NeRF 360 [5], a view synthesis dataset and on Tanks and Temples [50], a reconstruction dataset. Comparing poses on these datasets is problematic, as our pseudo ground truth is estimated rather than measured. For example, an approach might be more accurate than COLMAP on individual scenes. Computing pose errors w.r.t. COLMAP would result in incorrect conclusions. Therefore, we gauge the pose quality in a self-supervised way, using novel view synthesis [100].
We let each method estimate the poses of all images of a scene. For evaluation, we split the images into training and test sets. We train a Nerfacto [93] model on the training set, and synthesize views for the test poses. We compare the synthesized images to the test images and report the difference as peak signal-to-noise ratio (PSNR). To ensure a fair comparison to NeRF-based competitors, we use these methods in the same way as the other SfM methods: we run them to pose all images, and train a Nerfacto model on top of their poses. This is to ensure that we solely compare the pose quality across methods, and not different capabilities in view synthesis. The quality of poses affects the PSNR numbers in two ways: Good training poses let the NeRF model fit a consistent scene representation. Good testing poses make sure that the synthesized images are aligned with the original image. We explain further details in the supplement which also includes additional, perceptual metrics. In spirit, our evaluation is similar to the Tanks and Temples benchmark which evaluates a derived scene mesh rather than camera poses. However, our evaluation can be applied to arbitrary datasets as it does not need ground truth geometry.
@@
 Pseudo Ground Truth  All Frames  200 Frames  50 Frames Frames  Kinect  COLMAP  COLMAP  DROID-SLAM  ACE0  KF+ACE0  BARF  NoPE-NeRF  ACE0  DUSt3R  ACE0 Fusion (default) (fast) [94] (ours) (ours) [55] [10] (ours) [101] (Ours) Chess 6k 19.6 23.6  23.5  19.3  23.3  23.0 12.8 12.6 22.7 18.9 19.2 Fire 4k 19.2 22.6  22.6  13.0  22.3  22.3 12.7 11.8 22.1 18.8 19.5 Heads 2k 17.0 18.8  18.9  17.6  18.8  19.1 10.7 11.8 19.9 18.4 21.3 Office 10k 18.9 21.4  21.6 failed  21.1  21.5 11.9 10.9 19.8 12.5 13.7 Pumpkin 6k 19.9 24.1  23.8  18.3  24.1  23.8 19.6 14.2 24.7 21.7 22.3 RedKitchen 12k 17.6 21.4  21.4  10.9  20.8  20.9 11.6 11.2 18.9 13.8 13.7 Stairs 3k 19.0 16.7  21.0  13.0  17.7  19.9 15.8 15.9 18.8 15.3 15.4 Average 18.7 21.2  21.8 0N/A  21.2  21.5 13.6 12.6 21.0 17.1 17.9 Avg. Time realtime 38h  13h  18min  1h  7min 8.5h 47h 27min 4min* 16min
@@
Discussion. We show results in Table 1. Of both pseudo ground truth versions, KinectFusion achieves lower PSNR numbers than COLMAP, presumably due to the lack of global optimization. COLMAP with fast parameters shows PSNR numbers similar to COLMAP with default parameters, on average. Both versions of running COLMAP take considerable time to reconstruct each scene. We note that COLMAP has been optimised for quality, rather than speed. Not all acceleration strategies from the feature-based SfM literature have been implemented in COLMAP, so presumably comparable quality can be obtained faster. DROID-SLAM [94] does not perform well on 7-Scenes and partially fails altogether, presumably due to the jumps between individual scans of each scene.
Our approach, ACE0, achieves a pose quality comparable to the COLMAP pseudo ground truth while reconstructing each scene in $\sim$1 hour despite the large number of images. We show qualitative examples in Figure 4 and in the supplement. We also demonstrate that ACE0 can swiftly optimize an initial set of approximate poses. When starting from KinectFusion poses, ACE0 increases PSNR significantly in less than 10 minutes per scene, see “KF+ACE0” in Table 1. In the supplement, we include a parameter study on 7-Scenes to show that ACE0 is robust to the choice of depth estimator. We also show the positive impact of pose refinement on the reconstruction quality as well as the reconstruction speedup due to our early stopping schedule.
For our learning-based competitors, we sub-sampled images due to their computational constraints, see right side of Table 1. Even using only 200 images, NoPe-NeRF [10] takes 2 days to fit a model and estimate poses. Despite these long processing times, we observe poor pose quality of BARF and NoPe-NeRF in terms of PSNR. BARF [55] requires pose initialisation. We provide identity poses since the scenes of 7Scenes are roughly forward-facing. Still, the camera motions are too complex for BARF to handle. NoPe-NeRF does not require pose initialisation but needs roughly sequential images, which we did provide. NoPe-NeRF relies upon successive images having similar poses to perform well, and struggles with the large jumps between the subsampled images.